The sum of two numbers is 16. If one number is three times another number, what is the larger number?

Answer by stanbon(75887)

The sum of two numbers is 16. If one number is three times another number, what is the larger number?
-------
Equations:
x + y = 16
x = 3y
-----
Substitute for "x" and solve for "Y".
-----
3y + y = 16
4y = 16
y = 4
----
Solve for "x":
x = 3y
x = 3*4 = 12
===================
Cheers,
Stan H.
======================